OpenStudy (anonymous):

Which of the following is the solution to the equation 25^(z - 4) = 125 ? Select one: A. z = 5.5 B. z = 3.5 C. z = -2.5 D. z = -0.5

OpenStudy (anonymous):

ok. 25 = 5^2 and 125 = 5^3 we have - (5^2)(z-4) = 5^3 2z-8 = 3 2z = 11 or z = 5.5

OpenStudy (anonymous):

Thanks for the help!